# Modeling Advection-Dominated Flows with Space-Local Reduced-Order Models

### v2

@@ Abstract
Reduced-order models (ROMs) are often used to accelerate the simulation of large physical systems. However, traditional ROM techniques, such as those based on proper orthogonal decomposition (POD), often struggle with advection-dominated flows due to the slow singular value decay. This results in high computational costs and potential instabilities.
This paper proposes a novel approach using space-local POD to address the challenges arising from the slow singular value decay. Instead of global basis functions, our method employs local basis functions that are applied across the domain, analogous to the finite element method, but with a data-driven basis. By dividing the domain into subdomains and applying the space-local POD, we achieve a representation that is sparse and that generalizes better outside the training regime. This allows the use of a larger number of basis functions compared to standard POD, without prohibitive computational costs. To ensure smoothness across subdomain boundaries, we introduce overlapping subdomains inspired by the partition of unity method.
Our approach is validated through simulations of the 1D and 2D advection equation. We demonstrate that using our space-local approach we obtain a ROM that generalizes better to flow conditions which are not part of the training data. In addition, we show that the constructed ROM inherits the energy conservation and non-linear stability properties from the full-order model. Finally, we find that using a space-local ROM allows for larger time steps.
@@ 1 Introduction
Simulating large physical systems is an ongoing challenge in the field of computational sciences. This especially becomes challenging when dealing with multiscale systems. Such systems are characterized by displaying interesting behavior at different scales of space/time. A prominent example and our main incentive for this work are turbulent flows described by the incompressible Navier-Stokes equations. Systems described by these equations feature the formation of turbulent eddies of a range of different sizes. The large difference in size between the largest and the smallest eddies gives rise to the multiscale nature of the problem. The problem with simulating such systems is that they require high resolution computational meshes to obtain accurate simulations, and small time steps. This places a huge burden on the available computational resources [1, 2].
@@
To make simulation of turbulent flows feasible one typically resorts to Reynolds averaged Navier-Stokes [3], large eddy simulation [4], and reduced-order models (ROMs)[5, 6]. Here we focus on the latter approach. In reduced-order modelling, data is used to speed up simulations. The data can both be obtained from simulations or experiments. The collected data is then used to identify the most important features of the flow. This is typically done by a proper orthogonal decomposition (POD)of the collected flow data. The resulting features are then used to construct a reduced basis. By projecting the fluid flow equations on this basis one obtains the widely used POD-Galerkin ROM [7, 8, 6]. However, two common issues with POD-Galerkin ROMs are their stability and their accuracy for convection-dominated systems. In [9] it is shown that the stability issue can be resolved by making sure that the energy-conserving property of the Navier-Stokes equations is still satisfied under Galerkin projection. For the incompressible Navier-Stokes equations, the only prerequisite for an energy-conserving ROM is that the discretization is structure-preserving, i.e. it is such that it inherits the energy conservation property from the continuous equations.
@@
However, the stability property derived in [9] does not guarantee accuracy – it is possible to have energy-stable simulations that are highly inaccurate. This issue can already be observed on academic test cases such as the linear advection equation. For example, in [5] the authors discuss the problem of a single traveling wave, through a periodic domain. Even though it is clear that a one dimensional representation of the system exists, it is not recovered by the POD algorithm. The result is a slow decay in singular values of the snapshot matrix. This means that many POD modes are required to accurately describe the flow. The slow decay in singular values is often related to the Kolmogorov $N$-width. This is a measure of how well the solution space of a partial differential equation (PDE) can be represented by a linear combination of $N$ basis functions [10, 11]. Advection-dominated flows are notorious for displaying a slow Kolmogorov $N$-width decay, requiring a large $N$ for accurate simulation. The resulting ROM is then expensive to evaluate [12, 13]. When not including sufficient number of basis functions inaccurate results are obtained, e.g. they contain oscillations [14, 15].
@@
Different approaches to deal with this problem have been suggested. One way of dealing with this is by taking a local in time approach [16]. In this approach the ROM switches basis for different time intervals. This means that the ROM can employ a smaller basis, since each basis is specialized to deal with a certain time interval. Switching between basis during simulation can also be done in a structure-preserving (energy-conserving) manner [17]. A related approach is updating the basis during the ROM simulation to be more robust to changes in simulation conditions [18]. In [19] a Petrov-Galerkin approach is suggested, which leads to more stable ROMs when using a small POD basis than the standard Galerkin approach. In [20] it is suggested to add a closure model to the ROM to account for a small POD basis. The closure model is then tasked with modeling the interaction between the part of the flow that is covered by the POD basis and the part which is removed. Another suggestion is to construct ROMs on nonlinear reduced subspaces instead of linear ones. In [21] a more accurate representation of the solution space is obtained using rational quadratic manifolds. Here the ROM construction was also performed in a structure-preserving (entropy-stable) manner, yielding stability of the ROM.
@@
Machine learning approaches have also gained traction for the construction of ROMs. In [22] a long short-term memory (LSTM) neural network is used instead of Galerkin projection, as the computational complexity of LSTM is more favorable than Galerkin projection-based ROMs. They also allow one to take larger time steps [23]. In [23] this approach is combined with an autoencoder to reduce the dimensionality of the system. It is shown that a much greater reduction in degrees of freedom (DOF) can be reached using this approach compared to the standard POD-Galerkin approach. In [10] an advection-aware autoencoder is suggested to reduced the dimensionality of this system. This is achieved by introducing an additional decoder which is trained to reconstruct a “shifted” version of the encoded snapshots. This shift can for example be a snapshot taken further in time. This forces the latent space of the autoencoder to become aware of the dominant advective features of the flow.
@@
In this work we focus on another approach to tackle the issue of ROM accuracy in advection-dominated flows. Namely, we propose to use the idea of a space-local POD to tackle the slow singular value decay of advection-dominated flows. The idea is as follows: instead of obtaining a set of global basis functions which span the entire domain, we obtain a set of space-local basis functions. These basis functions are repeated across the entire domain, similarly to how a finite element basis covers the domain [24]. Furthermore, they are only nonzero on their designated subdomain. The advantage of this is that the resulting Galerkin projected operators are sparse. This makes them much cheaper to evaluate when simulating the ROM. For example, in [25] a speedup of up to 1.5 orders of magnitude is reported with respect to a standard POD-Galerkin ROM. Our approach differs from the one presented in [25]: In our approach, the solution data in each of the subdomains is treated with the same local basis in our local POD, instead of obtaining a different local POD basis for each subdomain as in [25]. This means that a feature observed in part of the domain, can also be represented on a different part of the domain, by the same basis. In this way, less data is required to obtain a POD basis that generalizes well. Note that this approach is designed specifically for problems where the dynamics are similar throughout the domain, such as a traveling wave. For problems where the dynamics are more variable throughout the domain the approach suggested by [25] is likely still superior, as it builds a specialized basis for each subdomain. Note that while our approach does not directly solve the slow Kolmogorov $N$-width decay (we still use linear approximations), the sparsity of the basis allows us to use a much larger number of basis functions. Sparsity and generalizability are therefore key features of our approach. In addition, we introduce another novel idea in this work: the use of overlapping subdomains (to avoid discontinuities at the subdomain boundaries). Furthermore, we show that our space-local ROMs satisfy energy conservation if the full-order model (FOM) does. In this way stability of the ROM is guaranteed.
@@
This paper is structured in the following way. In section 2 we introduce the FOM, namely a central difference discretization of the 1D advection equation. The central difference discretization ensures the scheme is energy conserving. In section 3 we introduce the POD approaches. We start with the standard POD approach and then introduce two space-local approaches, with and without overlapping subdomains. In section 4 we use Galerkin projection to project the FOM onto the POD basis and show that the resulting ROMs satisfy energy conservation. Finally, in section 5 we evaluate the different POD-based approaches in numerical experiments based on a set of different metrics. In addition, we evaluate the performance of the ROMs on a case where we extrapolate beyond the data used to construct the POD basis. Furthermore, we investigate the energy conserving properties of the ROMs. To conclude this section we apply the introduced methodologies to a 2D advection equation test case and evaluate the computational efficiency, among other metrics. Finally, in section 6, we present our main conclusions and suggest future research topics.
@@ 2.1 Advection equation
In this work we are concerned with developing a reduced-order model for the linear advection equation in both 1D and 2D. This equation is chosen as it exhibits similar difficulties as the incompressible Navier-Stokes equations when applying model reduction. For the sake of simplicity we discuss the properties of the system for 1D, but the ideas easily carry over to 2D. To start off, we consider a scalar solution $u(x,t)$ to the linear advection equation
@@
with initial condition $u(x,0)=u_{0}(x)$ and constant $c$. We employ a periodic domain $\Omega$ such that we do not require boundary conditions (BCs). An important property of this equation is that the total energy of the system
@@
$(a(x),b(x)):=\int_{\Omega}a(x)b(x)\text{d}\Omega.$ (3)
This can easily be shown using the product rule of differentiation:
$\frac{\text{d}E}{\text{d}t}=\frac{1}{2}\frac{\text{d}}{\text{d}t}(u,u)=(u,\frac{\partial u}{\partial t})=-c(u,\frac{\partial u}{\partial x})=c(\frac{\partial u}{\partial x},u)=0.$ (4)
In the final step we carried out integration-by-parts and used the fact that the boundary term cancels on periodic domains. The energy-conserving property of this equation will be mimicked by the discretization and by the ROMs developed in this work. This leads to unconditionally stable methods, as in [9].
@@ 2.2 Finite difference discretization
Although equation (1) can be easily solved exactly, for more complex equations this is not the case. In general, we approximate the solution by representing $u(x,t)$ on a grid. In this case, we employ a uniform grid with grid-spacing $h=\frac{|\Omega|}{N}$ such that $\text{u}_{i}(t)\approx u(x_{i},t)$. The approximated solution is contained within the state vector $\mathbf{u}(t)\in\mathbb{R}^{N}$. To approximate the spatial derivative we use a central difference approximation:
@@
where the linear operator $\mathbf{D}\in\mathbb{R}^{N\times N}$ is skew-symmetric and encodes the stencil in (5). For the time integration we use a classic Runge-Kutta 4 (RK4) scheme [26]. This time integration scheme induces a small energy conservation error, which will be negligible in our test cases. Alternatively, one can use the implicit midpoint method for exact energy conservation in time [27]. Equation (6) will be regarded as the full-order model (FOM). We note that other discretization techniques such as FEM can also be employed to derive a FOM, and our framework in section 3.
@@ 2.3 Energy conservation of the FOM
It is well known that this FOM mimicks the energy-conservation property (4) in a discrete setting. This can be shown by defining the discretized energy as
$E_{h}:=\frac{h}{2}\mathbf{u}^{T}\mathbf{u}.$ (7)
Using the product rule we obtain the following evolution equation for the energy
@@
In the final step we used the fact that $\mathbf{D}$ is skew-symmetric, i.e. $\mathbf{D}=-\mathbf{D}^{T}$, to see that the energy is conserved using this stencil. This yields both stability and consistency with the continuous equation.
@@ 3 Local and global POD
In order to reduce the computational cost of solving the FOM we construct a ROM. For this purpose we use simulation data to construct a data-driven basis. This basis is obtained through a POD of the simulation data [5, 6]. The most common approach is to employ a global POD basis, defined over the entire simulation domain, similar to a Fourier basis. An alternative is a local basis, as proposed in [25]. In this section we discuss both approaches. We will present our version of the space-local POD framework for finite difference discretizations, see (5). However, the ideas could potentially also be applied to different discretization techniques by first projecting onto a finite difference basis, as discussed in section 3.2. Our space-local approach has some key differences to the one presented in [25], which will be highlighted.
@@ 3.1 Global POD
Here $\psi_{i}$ can represent for example a Fourier basis or a localized box function in the case of the presented finite difference discretization. In the finite difference case we have $\mathbf{c}(t)=\mathbf{u}(t)$ and
@@
For finite element discretizations one could use e.g. Löwdin orthogonalization to express the solution in terms of an orthogonal basis [28].
@@
where $s$ is the number of snapshots. We decompose this snapshot matrix using a singular value decomposition (SVD) [29]
@@
$\boldsymbol{\Phi}=\operatorname*{arg\,min}_{\mathbf{M}\in\mathbb{R}^{N\times r}}||\mathbf{X}-\mathbf{M}\mathbf{M}^{T}\mathbf{X}||^{2}_{F},$ (13)
under the orthonormality constraint $\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}=\mathbf{I}$ [7, 8, 5, 6]. We can project the coefficient vector onto this basis as follows:
@@
As the obtained basis functions span the entire domain we refer to this approach as global proper orthogonal decomposition (G-POD).

### v3

@@ Abstract
Reduced-order models (ROMs) are often used to accelerate the simulation of large physical systems. However, traditional ROM techniques, such as proper orthogonal decomposition (POD)-based methods, often struggle with advection-dominated flows due to the slow decay of singular values. This results in high computational costs and potential instabilities.
This paper proposes a novel approach using space-local POD to address the challenges arising from the slow singular value decay. Instead of global basis functions, our method employs local basis functions that are applied across the domain, analogous to the finite element method, but with a data-driven basis. By dividing the domain into subdomains and applying the space-local POD, we obtain a sparse representation that generalizes better outside the training regime. This allows the use of a larger number of basis functions compared to standard POD, without prohibitive computational costs. To ensure smoothness across subdomain boundaries, we introduce overlapping subdomains inspired by the partition of unity method.
Our approach is validated through simulations of the 1D and 2D advection equation. We demonstrate that using our space-local approach, we obtain a ROM that generalizes better to flow conditions not included in the training data. In addition, we show that the constructed ROM inherits the energy conservation and non-linear stability properties from the full-order model. Finally, we find that using a space-local ROM allows for larger time steps.
@@ 1 Introduction
Simulating large physical systems is an ongoing challenge in the field of computational sciences. This especially becomes challenging when dealing with multiscale systems. Such systems exhibit interesting behavior at various spatial and temporal scales. A prominent example and our main incentive for this work are turbulent flows described by the incompressible Navier-Stokes equations. Systems described by these equations feature the formation of turbulent eddies of a range of different sizes. The significant difference in size between the largest and the smallest eddies gives rise to the multiscale nature of the problem. The problem with simulating such systems is that they require high-resolution computational meshes to obtain accurate simulations and small time steps. This places a significant burden on the available computational resources [sasaki2002navier, agdestein2024discretizefirstfilternext].
@@
To make simulation of turbulent flows feasible, one typically resorts to Reynolds averaged Navier-Stokes [RANS], large eddy simulation [sagaut2006large], and reduced-order models (ROMs) [brunton2019data, projection_ROMs_survey]. Here we focus on the latter approach. In reduced-order modelling, data is used to speed up simulations. The data can be obtained from simulations or experiments. The collected data is then used to identify the most critical features of the flow. This is typically done by a proper orthogonal decomposition (POD) of the collected flow data. The resulting features are then used to construct a reduced basis. By projecting the fluid flow equations on this basis one obtains the widely used POD-Galerkin ROM [POD1, POD2, projection_ROMs_survey]. However, two common issues with POD-Galerkin ROMs are their stability and their accuracy for convection-dominated systems. In [podbenjamin], it is shown that the stability issue can be resolved by making sure that the energy-conserving property of the Navier-Stokes equations is still satisfied under Galerkin projection. For the incompressible Navier-Stokes equations, the only prerequisite for an energy-conserving ROM is that the discretization is structure-preserving, i.e. it is such that it inherits the energy conservation property from the continuous equations.
@@
However, the stability property derived in [podbenjamin] does not guarantee accuracy; it is possible to have energy-stable simulations that are highly inaccurate. This issue can already be observed on academic test cases such as the linear advection equation. For example, in [brunton2019data], the authors discuss the problem of a single traveling wave through a periodic domain. Even though it is clear that a one-dimensional representation of the system exists, it is not recovered by the POD algorithm. The result is a slow decay in singular values of the snapshot matrix. This means that many POD modes are required to accurately describe the flow. The slow decay in singular values is often related to the Kolmogorov $N$-width. This is a measure of how well the solution space of a partial differential equation (PDE) can be represented by a linear combination of $N$ basis functions [advection_autoencoder_and_n_width, n_width]. Advection-dominated flows are notorious for displaying a slow Kolmogorov $N$-width decay, requiring a large $N$ for accurate simulation. The resulting ROM is then expensive to evaluate [expensive_1, expensive_2]. When not including a sufficient number of basis functions, inaccurate results are obtained, e.g., they contain oscillations [oscillations_1, oscillations_2].
@@
Different approaches to deal with this problem have been suggested. One way of dealing with this is by taking a local in time approach [time_local_farhat]. In this approach, the ROM switches basis for different time intervals. This means that the ROM can employ a smaller basis, since each basis is specialized to deal with a specific time interval. Switching between bases during simulation can also be done in a structure-preserving (energy-conserving) manner [KLEIN2024112697]. A related approach is to update the basis during the ROM simulation to make it more robust to changes in simulation conditions [adaptive_basis]. In [Grimberg], a Petrov-Galerkin approach is suggested, which leads to more stable ROMs when using a small POD basis than the standard Galerkin approach. In [ROM_closure], it is suggested to add a closure model to the ROM to account for a small POD basis. The closure model is then tasked with modeling the interaction between the part of the flow that is covered by the POD basis and the part that is removed. Another suggestion is to construct ROMs on non-linear reduced subspaces instead of linear ones. In [klein2024entropystablemodelreductiononedimensional], a more accurate representation of the solution space is obtained using rational quadratic manifolds. Here, the ROM construction was also performed in a structure-preserving (entropy-stable) manner, yielding stability of the ROM.
@@
Machine learning approaches have also gained traction for the construction of ROMs. In [POD_LSTM], a long short-term memory (LSTM) neural network is used instead of Galerkin projection, as the computational complexity of LSTM is more favorable than Galerkin projection-based ROMs. They also allow one to take larger time steps [MUCKE2021101408]. In [MUCKE2021101408], this approach is combined with an autoencoder to reduce the dimensionality of the system. It is demonstrated that a significantly greater reduction in degrees of freedom (DOF) can be achieved using this approach compared to the standard POD-Galerkin approach. In [advection_autoencoder_and_n_width], an advection-aware autoencoder is suggested to reduce the dimensionality of this system. This is achieved by introducing an additional decoder which is trained to reconstruct a “shifted” version of the encoded snapshots. This shift can, for example, be a snapshot taken at a later time. This forces the latent space of the autoencoder to become aware of the dominant advective features of the flow.
@@
In this work, we focus on an alternative approach to address the issue of ROM accuracy in advection-dominated flows. Namely, we propose to use the idea of a space-local POD to tackle the slow singular value decay of advection-dominated flows. The idea is as follows: instead of obtaining a set of global basis functions that span the entire domain, we obtain a set of space-local basis functions. These basis functions are repeated across the entire domain, similarly to how a finite element basis covers the domain [FEM_book]. Furthermore, they are only nonzero on their designated subdomain. The advantage of this is that the resulting Galerkin projected operators are sparse. This makes them much cheaper to evaluate when simulating the ROM. For example, in [farhat_local_POD] a speedup of up to 1.5 orders of magnitude is reported with respect to a standard POD-Galerkin ROM. Our approach differs from the one presented in [farhat_local_POD]. In our approach, the solution data in each subdomain is treated with the same local POD basis, rather than obtaining a different local POD basis for each subdomain. This approach ensures that a feature observed in part of the domain can also be represented in a different part of the domain using the same basis. In this way, less data is required to obtain a POD basis that generalizes well. Note that this approach is designed specifically for problems where the dynamics are similar throughout the domain, such as a traveling wave. For problems where the dynamics are more variable throughout the domain, the approach suggested by [farhat_local_POD] is likely still superior, as it builds a specialized basis for each subdomain. For diffusion-dominated problems the standard space-global POD approach is likely still superior due to the rapid Kolmogorov $N$-width decay [Quarteroni2015ROM]. Note that while our approach does not directly solve the slow Kolmogorov $N$-width decay (we still use linear approximations), the sparsity of the basis allows us to use a much larger number of basis functions. Sparsity and generalizability are therefore key features of our approach. Furthermore, the authors note this approach is similar to the methodology described in [CHUNG2024DDROM] where different types of subdomains were specified, each with a shared POD basis. These subdomains were then used as building blocks to extrapolate the POD basis obtained from small spatial domains to larger domains. Our work differs in the fact that it focuses on temporal extrapolation for time-dependent PDEs, as opposed to steady-state problems. In addition, we introduce another novel idea in this work: the use of overlapping subdomains (to avoid discontinuities at the subdomain boundaries). Lastly, we show that our space-local ROMs satisfy energy conservation if the full-order model (FOM) does. In this way, stability of the ROM is guaranteed.
@@
This paper is structured in the following way. In Section 2 we introduce the FOM, namely a central difference discretization of the 1D advection equation. The central difference discretization ensures the scheme is energy conserving. In Section 3 we introduce the POD approaches. We begin with the standard POD approach and then introduce two space-local approaches, one with and one without overlapping subdomains. In Section 4 we use Galerkin projection to project the FOM onto the POD basis and show that the resulting ROMs satisfy energy conservation. Finally, in Section 5, we evaluate the different POD-based approaches in numerical experiments using a set of different metrics. In addition, we assess the performance of the ROMs on a case where we extrapolate beyond the data used to construct the POD basis. Furthermore, we investigate the energy-conserving properties of the ROMs. To conclude this section, we apply the introduced methodologies to a 2D advection equation test case and evaluate the computational efficiency, among other metrics. Finally, in Section 6, we present our main findings and suggest future research topics.
@@ 2.1 Advection equation
In this work, we focus on developing a reduced-order model for the linear advection equation in both one-dimensional (1D) and two-dimensional (2D) settings. This equation is chosen as it exhibits similar difficulties as the incompressible Navier-Stokes equations when applying model reduction. For the sake of simplicity, we discuss the properties of the system for 1D, but the ideas easily carry over to 2D. To start, we consider a scalar solution $u(x,t)$ to the linear advection equation
@@
with initial condition $u(x,0)=u_{0}(x)$ and constant $c$. In this work we stick to periodic boundary conditions (BCs). An important property of this equation is that the total energy of the system
@@
$(a(x),b(x)):=\int_{\Omega}a(x)b(x)\text{d}\Omega$ (3)
on the spatial domain $\Omega$. This can easily be shown using the product rule of differentiation:
$\frac{\text{d}E}{\text{d}t}=\frac{1}{2}\frac{\text{d}}{\text{d}t}(u,u)=\left(u,\frac{\partial u}{\partial t}\right)=-c\left(u,\frac{\partial u}{\partial x}\right)=c\left(\frac{\partial u}{\partial x},u\right)=0.$ (4)
In the final step, we carried out integration by parts and used the fact that the boundary term cancels on periodic domains. The energy-conserving property of this equation will be mimicked by the discretization and by the ROMs developed in this work. This leads to unconditionally stable methods, as in [podbenjamin].
@@ 2.2 Finite difference discretization
Although equation (1) can be easily solved exactly, for more complex equations, this is not the case. In general, we approximate the solution by representing $u(x,t)$ on a grid. In this case, we employ a uniform grid with grid spacing $h=\frac{|\Omega|}{N}$ such that $\text{u}_{i}(t)\approx u(x_{i},t)$. The approximated solution is contained within the state vector $\mathbf{u}(t)\in\mathbb{R}^{N}$. To approximate the spatial derivative, we use a central difference approximation:
@@
where the linear operator $\mathbf{D}\in\mathbb{R}^{N\times N}$ is skew-symmetric and encodes the stencil in (5). For the time integration, we use a classic Runge-Kutta 4 (RK4) scheme [RK4_Butcher:2007]. This time integration scheme introduces a small energy conservation error, which is negligible in our test cases. Alternatively, one can use the implicit midpoint method for exact energy conservation in time [SANDERSE2013]. Equation (6) will be regarded as the full-order model (FOM). We note that other discretization techniques, such as the finite element method, can also be employed to derive a FOM; our framework in Section 3 is also applicable in this context.
@@ 2.3 Energy conservation of the FOM
It is well known that this FOM mimics the energy-conservation property (4) in a discrete setting. This can be shown by defining the discretized energy as
$E_{h}:=\frac{h}{2}\mathbf{u}^{T}\mathbf{u},$ (7)
where the inclusion of $h$ is such that this definition discretely represents the inner product in (2).
Using the product rule, we obtain the following evolution equation for the energy
@@
In the final step, we used the fact that $\mathbf{D}$ is skew-symmetric, i.e., $\mathbf{D}=-\mathbf{D}^{T}$, to see that the energy is conserved using this stencil. This yields both stability and consistency with the continuous equation.
@@ 3 Local and global POD
To reduce the computational cost of solving the FOM, we construct a ROM. For this purpose, we use simulation data to build a data-driven basis. This basis is obtained through a POD of the simulation data [brunton2019data, projection_ROMs_survey]. The most common approach is to employ a global POD basis, defined over the entire simulation domain, similar to a Fourier basis. An alternative is a local basis, as proposed in [farhat_local_POD, CHUNG2024DDROM]. In this section, we discuss both approaches. We will present our version of the space-local POD framework for finite difference discretizations, see (5). However, the ideas can also be applied to different discretization techniques, as shown in [CHUNG2024DDROM]. Our space-local approach has some key differences from the one presented in [farhat_local_POD], which will be highlighted.
@@ 3.1 Global POD
Here $\psi_{i}$ can represent, for example, a Fourier basis or a localized box function in the case of the presented finite difference discretization. In the finite difference case we have $\mathbf{c}(t)=\mathbf{u}(t)$ and
@@
For finite element discretizations, one could use, e.g., Löwdin orthogonalization to express the solution in terms of an orthogonal basis [aiken1980lowdin].
@@
where $s$ is the number of snapshots. We decompose this snapshot matrix using a singular value decomposition (SVD) [SVD_complexity]
@@
[…]
under the orthonormality constraint $\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}=\mathbf{I}$ [POD1, POD2, brunton2019data, projection_ROMs_survey]. We can project the coefficient vector onto this basis as follows:
@@
As the obtained basis functions span the entire domain, we refer to this approach as space-global proper orthogonal decomposition (G-POD).